OpenStudy (i_love_my_nieces):

which ordered pair is NOT an solution? y=-4x - 9 ? (-9, 27) (3, -21) (-6, 15) (5, -28)

OpenStudy (i_love_my_nieces):

@mathstudent55

OpenStudy (solomonzelman):

plug your solutions in and see which ones work :)

OpenStudy (i_love_my_nieces):

@dan815

OpenStudy (mathstudent55):

You need to try them one by one until you find a choice that does not work. y = -4x - 9 Let's try choice C. (-6, 15) Replace x with -6 and y with 15 in the given equation. If the equation is true, then the ordered pair works. 15 = -4(-6) - 9 15 = 24 - 9 15 = 15 This is a true statement, so choice C. is a solution. Do the same to choices A, B, and D, and find which one does not work.

OpenStudy (i_love_my_nieces):

Thank you

OpenStudy (mathstudent55):

wlcm